Suppose you know a company's stock currently sells for $60 per share and the required return on the stock is 9 percent. You also know that the total return on the stock is evenly divided between a capital gains yield and a dividend yield.

  

If it's the company's policy to always maintain a constant growth rate in its dividends, what is the current dividend per share?

Please be specific and provide formulas
Answer #1

Required return = dividend yield + capital gains yield

Since required return is evenly divided between two, dividend yield is 4.5% and capital gains yield to 5%

Dividend yield = D1 / share price

0.045 = D1 / 60

D1 = 2.7

Current dividend = D1 / (1 + g)

g is capital gains yield

Current dividend = 2.7 / (1 + 0.045)

Current dividend = 2.58